# Bayesian inference and cure rate modeling for event history data

### v1

@@ Abstract
Estimating model parameters of a general family of cure models is always a challenging task mainly due to flatness and multimodality of the likelihood function. In this work, we propose a fully Bayesian approach in order to overcome these issues. Posterior inference is carried out by constructing a Metropolis-coupled Markov chain Monte Carlo (MCMC) sampler, which combines Gibbs sampling for the latent cure indicators and Metropolis-Hastings steps with Langevin diffusion dynamics for parameter updates. The main MCMC algorithm is embedded within a parallel tempering scheme by considering heated versions of the target posterior distribution. It is demonstrated via simulations that the proposed algorithm freely explores the multimodal posterior distribution and produces robust point estimates, while it outperforms maximum likelihood estimation via the Expectation-Maximization algorithm. A by-product of our Bayesian implementation is to control the False Discovery Rate when classifying items as cured or not. Finally, the proposed method is illustrated in a real dataset which refers to recidivism for offenders released from prison; the event of interest is whether the offender was re-incarcerated after probation or not.
@@ 1 Introduction
where $\pi\in[0,1]$ is the probability of being cured, known as cure rate (incidence), and $S(t|I=1)$ is the survival function of susceptibles (latency); due to the definition of cured subjects, we assume that $S(t|I=0)=1$, for every $t$. In fact, the story of cure modeling started with model (1), the mixture cure model, and the seminal works by Boag [11, 12, 13] which rigorously treated the idea of the existence of cured patients, among those who get a cancer treatment. However, real-life applications make researchers consider different ways of splitting the population into groups, such the one deduced by a competing cause scenario (the roots of this approach can be found in 14, 15, 16, although some earlier works of similar nature could also be met in 17, 18); specifically, suppose that there is a number of competing causes, $M$, with each cause being able (at least, partially) to deliver the event of interest. Then, the population survival function is now given by
@@
Although a great part of the literature refers to the study of the mixture cure model, also known as split population model [6, Ch. 5], or limited-failure population model [19, 20], the research interest, at least in the last two decades, is also heavily focused on the competing cause approach and its extensions. Typically, the time needed for the $j$th competing cause to being activated (promotion or progression time) is a random variable, $W_{j}$, and given $M$, $W_{j},j=1,2,\ldots$ are assumed to be independent and identically distributed, and independent of $M$, with a common distribution function $F(t)=1-S(t)$. Then, if the first activation suffices for delivering the event of interest, i.e., $T=\min\{W_{0},\ldots,W_{M}\}$, with $P(W_{0}=\infty)=1$, model (2) becomes
@@
where $\varphi(z)$ is the probability generating function of $M$ (e.g., 21). Other approaches generalize the above classical competing cause scenario, by asking, for example, the activation of $r$ out of $M$ causes, for facing the event of interest, with $r\geq 1$ (e.g., 22; some generalizations could also be found in 21, 23, 24, 25). Obviously, the stochastic properties of $M$ play a critical role to the competing cause cure model, and one of the fundamental points for selecting a distribution for $M$ (apart from being non-negative integer-valued random variable, with mass at zero) is whether this distribution is flexible enough, to deal with many different real-life settings. The case where $M$ follows a Poisson distribution, i.e., $S_{p}(t)=\exp\{-\vartheta F(t)\}$, with $\vartheta>0$ being the parameter of the Poisson distribution, known as bounded cumulative hazard or promotion time cure model, is one of the key references in the literature. Having said that, the research interest often focuses on proposing models which contain as special cases the two most studied cure models: the mixture cure model and the promotion time model. This is the case, for example, for the model derived by using the Box-Cox transformation on the population survival function, which is of the form (26)
@@
with $\vartheta:\rightarrow\delta/(1+\delta\gamma)$, and $\delta>0$ (see also 27, 28, 29, 30, 31, 32). Note that both the promotion time ($\gamma\rightarrow 0$) and mixture cure model ($\gamma=1$) are special cases of this family; $\vartheta$ was given by the above form due to the restriction $\gamma\vartheta\in[0,1]$. To mention that this model was estimated within a Bayesian framework in [26] and posterior inference was performed using Adaptive Rejection Metropolis sampling within the Gibbs sampler [33]. Of similar nature are the models derived by assuming a negative binomial distribution for the number of competing causes (e.g., 34, 35, 36, 37, 38, 39, 40, 41), or a Poisson distribution with random parameter (e.g., 42) or a generalized discrete Linnik distribution (43).
@@
The rest of the paper is organized as follows. Section 2 presents the main ingredients of our Bayesian approach, that is, the complete likelihood after augmenting the model with latent status indicators, as well as the prior assumptions. Section 3 introduces the MCMC sampling schemes for performing posterior inference. Control of the False Discovery Rate is discussed in Section 4. Illustrations of the proposed methodology are given in Section 5 using simulated datasets and comparing against the EM algorithm (Section 5.1); also, a real-life application on recidivism for offenders released from prison (parole/special sentence, work release, or discharge) is presented in Section 5.2. Further details of the implementation and additional results on the simulation study are included in the Appendix.
@@ 2 Bayesian model
The estimation of the distribution of susceptibles and the recognition of the cured subjects, are conventionally the most challenging tasks in cure modeling. The study of such populations is commonly characterized by the existence of censored data; this is because the cured subjects do not face the event of interest and therefore, no matter what censoring scheme (research design) has been followed up, their time-to-event are censored and their cure status is unknown (nevertheless, the cure status is being treated as partially known, in some approaches; e.g., 49, 50, 51, 52, 53, 54).
In our approach, we assume that the time-to-event is subject to non-informative random right censoring. Under the classical competing cause scenario and model (4), we further assume that the promotion time $W_{j}$ follows a Weibull distribution; however, as can also be seen by the presentation of our estimation approach, this assumption can effortlessly be relaxed. This is also the case about the role of covariates; although we assume that the set of covariates affects only parameter $\vartheta$, through an exponential link function, the same or another set of covariates (under conditions), could also affect the rest parameters. Furthermore, it is known that there are some identifiability issues with the model under study (see also 44). For example, under (3), we can find two different sets of parameters with exactly the same survival function, given that no covariates have been used in the model; also, when $\gamma=-e$, $\lambda=1$ and the distribution of the promotion time is a Weibull distribution, then a constant cannot be added to the exponential link function of $\vartheta$, in case of using covariates. Here, we use the parametrization (4) and further assuming the existence of a continuous covariate with non-negligible effect on $\vartheta$ through the exponential link function, the identifiability issues are solved.
@@ 2.2 Prior assumptions
$\vartheta=\vartheta({\boldsymbol{x}}_{i})=\exp\{{\boldsymbol{\beta}}{\boldsymbol{x}}_{i}^{\prime}\}=\exp\{\beta_{0}+\beta_{1}x_{1}+\ldots+\beta_{k}x_{k}\},$
@@
$\gamma$ $\sim f_{\gamma}(a_{\gamma},b_{\gamma})$ (6)
@@
where $a_{\gamma}>0$, $b_{\gamma}>0$, $a_{1}>0$, $a_{2}>0$, $b_{1}>0$, $b_{2}>0$, $a_{\lambda}>0$, $b_{\lambda}>0$, $\boldsymbol{\mu}\in\mathbb{R}^{k+1}$, $\Sigma\in\mathcal{M}_{+}$, with $\mathcal{M}_{+}$ denoting the space of $(k+1)\times(k+1)$ symmetric positive definite matrices ($G(\cdot,\cdot),\mathcal{IG}(\cdot,\cdot)$ and $\mathcal{N}_{p}(\cdot,\cdot)$, denote the Gamma (shape, rate), inverse Gamma and $p$th dimensional normal distribution, respectively). The prior distribution $f_{\gamma}(a_{\gamma},b_{\gamma})$ of $\gamma$ denotes an equally weighted mixture of a $\mathcal{G}(a_{\gamma},b_{\gamma})$ distribution and its reflection towards 0, that is,
$\gamma=\begin{cases}\hfill Z,\quad&\mbox{with probability }1/2\\ -Z,\quad&\mbox{with probability }1/2\end{cases}$
where $Z\sim\mathcal{G}(a_{\gamma},b_{\gamma})$.
The prior assumptions stated above are based on adopting continuous distributions on the parameter space. Although this seems a reasonable approach, it is not always the case in cure modeling theory; for example, we refer the reader to [26] where a discrete prior distribution was assigned to $\gamma$, under the Box-Cox transformation cure model, in order to avoid numerical problems. The main reason for choosing prior (6) is that it allows greater flexibility compared to a unimodal distribution (e.g. a normal distribution centered at zero). For the remaining parameters, we use absolutely continuous prior distributions on their support, something that can be achieved by inverse Gamma ($\mathcal{IG}$) distributions for the positive parameters and normal distributions for the regression coefficients ($\boldsymbol{\beta}$). In our applications we consider two distinct sets of hyperparameter values, corresponding to “vague” and “regularized” setups. The reader is referred to Appendix A for more details. Figure 1 illustrates the previous hierarchical cure model as a directed acyclic graph, where circles are used to denote unobserved (partially, at least) random variables and squares represent fixed/observed quantities.
@@ 3 MCMC sampling schemes
$\pi(\boldsymbol{\theta},\boldsymbol{I}|\boldsymbol{\Delta},\boldsymbol{y},\boldsymbol{x})\propto$ $f(\boldsymbol{y},\boldsymbol{I}|\boldsymbol{\Delta},\boldsymbol{\theta},\boldsymbol{x})\pi(\boldsymbol{\theta})$
@@ 3.1 Metropolis-Hastings and MALA updates of parameters
More specifically, the first part of our MCMC sampler consists of standard Metropolis-Hastings updates. In brief, we have used a normal proposal distribution for $\gamma$, log-normal distributions for $\alpha_{1}$, $\alpha_{2}$ and $\lambda$ and a multivariate normal proposal for $\boldsymbol{\beta}$ (see Appendix B.1 for detailed descriptions of these moves).
@@
where $\boldsymbol{\varepsilon}\sim\mathcal{N}(\boldsymbol{0},\mathcal{I}_{p+4})$ and $\nabla\log\pi^{h}(\boldsymbol{\theta}|\boldsymbol{\Delta},\boldsymbol{y},\boldsymbol{x},\boldsymbol{I})$ denotes the gradient vector of the logarithm of the full conditional heated posterior distribution of $\boldsymbol{\theta}$ ($\mathcal{I}_{p+4}$ stands for the $(p+4)\times(p+4)$ identity matrix; see Appendices B.2 and C for a detailed description of the implementation).
@@ 3.2 Gibbs update for latent cured status indicators
    Generate $\tilde{\gamma}$ according to (14) and set $\gamma^{(t)}=\tilde{\gamma}$ with probability $A_{1,h}(\gamma^{(t-1)},\tilde{\gamma})$ in Equation (15), otherwise set $\gamma^{(t)}=\gamma^{(t-1)}$.
@@
    Generate $\tilde{\lambda}$ according to (16) and set $\lambda^{(t)}=\tilde{\lambda}$ with probability $A_{2,h}(\lambda^{(t-1)},\tilde{\lambda})$ in Equation (17), otherwise set $\lambda^{(t)}=\lambda^{(t-1)}$.
@@
    Generate $\tilde{\alpha}_{1}$ according to (18) and set $\alpha_{1}^{(t)}=\tilde{\alpha}_{1}$ with probability $A_{3,h}(\alpha_{1}^{(t-1)},\tilde{\alpha}_{1})$ in Equation (19), otherwise set $\alpha_{1}^{(t)}=\alpha_{1}^{(t-1)}$.
@@
    Generate $\tilde{\alpha}_{2}$ according to (20) and set $\alpha_{2}^{(t)}=\tilde{\alpha}_{2}$ with probability $A_{4,h}(\alpha_{2}^{(t-1)},\tilde{\alpha}_{2})$ in Equation (21), otherwise set $\alpha_{2}^{(t)}=\alpha_{2}^{(t-1)}$.
@@
    Generate $\tilde{\boldsymbol{\beta}}$ according to (22) and set $\boldsymbol{\beta}^{(t)}=\tilde{\boldsymbol{\beta}}$ with probability $A_{5,h}(\boldsymbol{\beta}^{(t-1)},\tilde{\boldsymbol{\beta}})$ in Equation (23), otherwise set $\boldsymbol{\beta}^{(t)}=\boldsymbol{\beta}^{(t-1)}$.
@@
             Generate $\tilde{\boldsymbol{\theta}}$ according to (10) and set $\boldsymbol{\theta}^{(t)}=\tilde{\boldsymbol{\theta}}$ with probability $A_{6,h}(\boldsymbol{\theta}^{(t-1)},\tilde{\boldsymbol{\theta}})$ in Equation (24), otherwise set $\boldsymbol{\theta}^{(t)}=\boldsymbol{\theta}^{(t-1)}$.

### v2

@@ Abstract
Estimating model parameters of a general family of cure models is always a challenging task mainly due to flatness and multimodality of the likelihood function. In this work, we propose a fully Bayesian approach in order to overcome these issues. Posterior inference is carried out by constructing a Metropolis-coupled Markov chain Monte Carlo (MCMC) sampler, which combines Gibbs sampling for the latent cure indicators and Metropolis-Hastings steps with Langevin diffusion dynamics for parameter updates. The main MCMC algorithm is embedded within a parallel tempering scheme by considering heated versions of the target posterior distribution. It is demonstrated that along the considered simulation study the proposed algorithm freely explores the multimodal posterior distribution and produces robust point estimates, while it outperforms maximum likelihood estimation via the Expectation-Maximization algorithm. A by-product of our Bayesian implementation is to control the False Discovery Rate when classifying items as cured or not. Finally, the proposed method is illustrated in a real dataset which refers to recidivism for offenders released from prison; the event of interest is whether the offender was re-incarcerated after probation or not.
@@ 1 Introduction
where $\pi\in[0,1]$ is the probability of being cured, known as cure rate (incidence), and $S(t|I=1)$ is the survival function of susceptibles (latency); due to the definition of cured subjects, we assume that $S(t|I=0)=1$, for every $t$. In fact, the story of cure modeling started with model (1), the mixture cure model, and the seminal works by Boag [11] which rigorously treated the idea of the existence of cured patients, among those who get a cancer treatment. However, real-life applications make researchers consider different ways of splitting the population into groups, such the one deduced by a competing cause scenario (the roots of this approach can be found in 12, 13, 14, although some earlier works of similar nature could also be met in 15, 16); specifically, suppose that there is a number of competing causes, $M$, with each cause being able (at least, partially) to deliver the event of interest. Then, the population survival function is now given by
@@
Although a great part of the literature refers to the study of the mixture cure model, also known as split population model [6, Ch. 5], or limited-failure population model [17], the research interest, at least in the last two decades, is also heavily focused on the competing cause approach and its extensions. Typically, the time needed for the $j$th competing cause to being activated (promotion or progression time) is a random variable, $W_{j}$, and given $M$, $W_{j},j=1,2,\ldots$ are assumed to be independent and identically distributed, and independent of $M$, with a common distribution function $F(t)=1-S(t)$. Then, if the first activation suffices for delivering the event of interest, i.e., $T=\min\{W_{0},\ldots,W_{M}\}$, with $P(W_{0}=\infty)=1$, model (2) becomes
@@
where $\varphi(z)$ is the probability generating function of $M$ (e.g., 18). Other approaches generalize the above classical competing cause scenario, by asking, for example, the activation of $r$ out of $M$ causes, for facing the event of interest, with $r\geq 1$ (e.g., 19; some generalizations could also be found in 18, 20, 21, 22). Obviously, the stochastic properties of $M$ play a critical role to the competing cause cure model, and one of the fundamental points for selecting a distribution for $M$ (apart from being non-negative integer-valued random variable, with mass at zero) is whether this distribution is flexible enough, to deal with many different real-life settings. The case where $M$ follows a Poisson distribution, i.e., $S_{P}(t)=\exp\{-\vartheta F(t)\}$, with $\vartheta>0$ being the parameter of the Poisson distribution, known as bounded cumulative hazard or promotion time cure model, is one of the key references in the literature. Having said that, the research interest often focuses on proposing models which contain as special cases the two most studied cure models: the mixture cure model and the promotion time model. This is the case, for example, for the model derived by using the Box-Cox transformation on the population survival function, which is of the form (23)
@@
with $\vartheta:\rightarrow\delta/(1+\delta\gamma)$, and $\delta>0$ (see also 24, 25, 26, 27, 28, 29). Note that both the promotion time ($\gamma\rightarrow 0$) and mixture cure model ($\gamma=1$) are special cases of this family; $\vartheta$ was given by the above form due to the restriction $\gamma\vartheta\in[0,1]$. To mention that this model was estimated within a Bayesian framework in [23] and posterior inference was performed using Adaptive Rejection Metropolis sampling within the Gibbs sampler [30]. Of similar nature are the models derived by assuming a negative binomial distribution for the number of competing causes (e.g., 31, 32, 33, 34, 35), or a Poisson distribution with random parameter (e.g., 36) or a generalized discrete Linnik distribution (37).
@@
The rest of the paper is organized as follows. Section 2 presents the main ingredients of our Bayesian approach, that is, the complete likelihood after augmenting the model with latent status indicators, as well as the prior assumptions. Section 3 introduces the MCMC sampling schemes for performing posterior inference. Control of the False Discovery Rate is discussed in Section 4. Illustrations of the proposed methodology are given in Section 5 using simulated datasets and comparing against the EM algorithm (Section 5.1); also, a real-life application on recidivism for offenders released from prison (parole/special sentence, work release, or discharge) is presented in Section 5.2. Further details of the implementation and additional results on the simulation study are included in Supplementary Material.
@@ 2 Bayesian model
The estimation of the distribution of susceptibles and the recognition of the cured subjects, are conventionally the most challenging tasks in cure modeling. The study of such populations is commonly characterized by the existence of censored data; this is because the cured subjects do not face the event of interest and therefore, no matter what censoring scheme (research design) has been followed up, their time-to-event are censored and their cure status is unknown (nevertheless, the cure status is being treated as partially known, in some approaches; e.g., 43, 44, 45, 46, 47, 48).
In our approach, we assume that the time-to-event is subject to non-informative random right censoring. Under the classical competing cause scenario and model (4), we further assume that the promotion time $W_{j}$ follows a Weibull distribution; however, as can also be seen by the presentation of our estimation approach, this assumption can effortlessly be relaxed. This is also the case about the role of covariates; although we assume that the set of covariates affects only parameter $\vartheta$, through an exponential link function, the same or another set of covariates (under conditions), could also affect the rest parameters. Furthermore, it is known that there are some identifiability issues with the model under study (see also 38). For example, under (3), we can find two different sets of parameters with exactly the same survival function, given that no covariates have been used in the model; also, when $\gamma=-e$, $\lambda=1$ and the distribution of the promotion time is a Weibull distribution, then a constant cannot be added to the exponential link function of $\vartheta$, in case of using covariates. Here, we use the parametrization (4) and further assuming the existence of a continuous covariate with non-negligible effect on $\vartheta$ through the exponential link function, the identifiability issues are solved.
@@ 2.2 Prior assumptions
$\vartheta_{i}:=\vartheta({\boldsymbol{x}}_{i})=\exp\{{\boldsymbol{\beta}}{\boldsymbol{x}}_{i}^{\prime}\}=\exp\{\beta_{0}+\beta_{1}x_{i1}+\ldots+\beta_{k}x_{ik}\},$
@@
$\gamma$ $\sim f_{\gamma}(a_{\gamma},b_{\gamma}):=\frac{b_{\gamma}^{a_{\gamma}}}{2\Gamma(a_{\gamma})}|\gamma|^{a_{\gamma}-1}\exp\{-b_{\gamma}|\gamma|\}\mathrm{I}_{\mathbb{R}^{\star}}(\gamma)$ (6)
@@
where $a_{\gamma}>0$, $b_{\gamma}>0$, $a_{1}>0$, $a_{2}>0$, $b_{1}>0$, $b_{2}>0$, $a_{\lambda}>0$, $b_{\lambda}>0$, $\boldsymbol{\mu}\in\mathbb{R}^{k+1}$, $\Sigma\in\mathcal{M}_{+}$, with $\mathcal{M}_{+}$ denoting the space of $(k+1)\times(k+1)$ symmetric positive definite matrices ($G(\cdot,\cdot),\mathcal{IG}(\cdot,\cdot)$ and $\mathcal{N}_{p}(\cdot,\cdot)$, denote the Gamma (shape, rate), inverse Gamma and $p$th dimensional normal distribution, respectively). Finally, $f_{\gamma}$ corresponds to the density of a random variable defined as $\gamma=ZU$ where $Z\sim\mathcal{G}(a_{\gamma},b_{\gamma})$ and $U\sim\mathcal{U}niform_{\{-1,1\}}$, with $Z$ and $U$ independent. Note that (6) reduces to a standard Laplace distribution when $a_{\gamma}=b_{\gamma}=1$ (a choice which will be considered in our simulations, among others).
The prior assumptions stated above are based on adopting continuous distributions on the parameter space. Although this seems a reasonable approach, it is not always the case in cure modeling theory; for example, we refer the reader to [23] where a discrete prior distribution was assigned to $\gamma$, under the Box-Cox transformation cure model, in order to avoid numerical problems. In our applications we consider two distinct sets of hyperparameter values, corresponding to “vague” and “regularized” setups. The reader is referred to Section A of supplementary material, for more details. Figure 1 illustrates the previous hierarchical cure model as a directed acyclic graph, where circles are used to denote unobserved (partially, at least) random variables and squares represent fixed/observed quantities.
@@ 3 MCMC sampling schemes
$\pi(\boldsymbol{\theta},\boldsymbol{I}|\boldsymbol{\Delta},\boldsymbol{y},\boldsymbol{x})\propto$ $f(\boldsymbol{y}|\boldsymbol{\theta},\boldsymbol{I},\boldsymbol{\Delta},\boldsymbol{x})\pi(\boldsymbol{I},\boldsymbol{\theta})$
$\propto$ $f(\boldsymbol{y}|\boldsymbol{\theta},\boldsymbol{I},\boldsymbol{\Delta},\boldsymbol{x})\pi(\boldsymbol{I}|\boldsymbol{\theta})\pi(\boldsymbol{\theta})$
$\propto$ $f(\boldsymbol{y},\boldsymbol{I}|\boldsymbol{\Delta},\boldsymbol{\theta},\boldsymbol{x})\pi(\boldsymbol{\theta})$ (7)
@@ 3.1 Metropolis-Hastings and MALA updates of parameters
More specifically, the first part of our MCMC sampler consists of standard Metropolis-Hastings updates. In brief, we have used a normal proposal distribution for $\gamma$, log-normal distributions for $\alpha_{1}$, $\alpha_{2}$ and $\lambda$ and a multivariate normal proposal for $\boldsymbol{\beta}$ (see Section B.1 of supplementary material for detailed descriptions of these moves).
@@
where $\boldsymbol{\varepsilon}\sim\mathcal{N}(\boldsymbol{0},\mathcal{I}_{p+4})$ and $\nabla\log\pi^{h}(\boldsymbol{\theta}|\boldsymbol{\Delta},\boldsymbol{y},\boldsymbol{x},\boldsymbol{I})$ denotes the gradient vector of the logarithm of the full conditional heated posterior distribution of $\boldsymbol{\theta}$ ($\mathcal{I}_{p+4}$ stands for the $(p+4)\times(p+4)$ identity matrix; see Sections B.2 and C of supplementary material for a detailed description of the implementation).
@@ 3.2 Gibbs update for latent cured status indicators
    Generate $\tilde{\gamma}$ according to (B.15) and set $\gamma^{(t)}=\tilde{\gamma}$ with probability $A_{1,h}(\gamma^{(t-1)},\tilde{\gamma})$ in Equation (B.16), otherwise set $\gamma^{(t)}=\gamma^{(t-1)}$.
@@
    Generate $\tilde{\lambda}$ according to (B.17) and set $\lambda^{(t)}=\tilde{\lambda}$ with probability $A_{2,h}(\lambda^{(t-1)},\tilde{\lambda})$ in Equation (B.18), otherwise set $\lambda^{(t)}=\lambda^{(t-1)}$.
@@
    Generate $\tilde{\alpha}_{1}$ according to (B.19) and set $\alpha_{1}^{(t)}=\tilde{\alpha}_{1}$ with probability $A_{3,h}(\alpha_{1}^{(t-1)},\tilde{\alpha}_{1})$ in Equation (B.20), otherwise set $\alpha_{1}^{(t)}=\alpha_{1}^{(t-1)}$.
@@
    Generate $\tilde{\alpha}_{2}$ according to (B.21) and set $\alpha_{2}^{(t)}=\tilde{\alpha}_{2}$ with probability $A_{4,h}(\alpha_{2}^{(t-1)},\tilde{\alpha}_{2})$ in Equation (B.22), otherwise set $\alpha_{2}^{(t)}=\alpha_{2}^{(t-1)}$.
@@
    Generate $\tilde{\boldsymbol{\beta}}$ according to (B.23) and set $\boldsymbol{\beta}^{(t)}=\tilde{\boldsymbol{\beta}}$ with probability $A_{5,h}(\boldsymbol{\beta}^{(t-1)},\tilde{\boldsymbol{\beta}})$ in Equation (B.24), otherwise set $\boldsymbol{\beta}^{(t)}=\boldsymbol{\beta}^{(t-1)}$.
@@
            Generate $\tilde{\boldsymbol{\theta}}$ according to (11) and set $\boldsymbol{\theta}^{(t)}=\tilde{\boldsymbol{\theta}}$ with probability $A_{6,h}(\boldsymbol{\theta}^{(t-1)},\tilde{\boldsymbol{\theta}})$ in Equation (B.25), otherwise set $\boldsymbol{\theta}^{(t)}=\boldsymbol{\theta}^{(t-1)}$.